# Well-posedness of a generalized Stokes operator on domains with cylindrical ends via layer-potentials

### 11 May 2026 v1

@@ 1.4. Overview of the existing literature
The analysis on non-compact manifolds poses more challenges than that the one on compact manifolds. A significant issue in this sense is the construction of the classes of pseudodifferential operators. In this regard, we refer to the contributions of Ammann, Lauter and Nistor [11, 12], Mazzeo and Melrose [84], Melrose [93], Melrose and Mendoza [95], Nistor [103], Nistor, Weinstein, and Xu [104] Schrohe [120, 121], Schulze [122, 123], Schulze and Wong [124], Shubin [128], to mention just a few among very many others.
Another important issue in the analysis on non-compact manifolds is finding the “right definition” of the relevant function spaces. In this paper, as in [66], we have defined our function spaces by using connections. See also [3, 4, 119]. Amann [4] and Ammann, Große, Nistor [6, 8] studied function spaces on manifolds with bounded geometry and obtained well-posedness results for the Laplace operator. Using the uniform Shapiro-Lopatinski conditions, Große and Nistor [49] obtained regularity results for elliptic boundary value problems on manifolds with bounded geometry. Lie manifolds are special manifolds with bounded geometry [10, 11], which allow the implementation of the layer potential methods. See also [101, 102].
@@
In [48], Große, Kohr, and Nistor have obtained the $L^{2}$-unique continuation property for the deformation operator $\operatorname{Def}\,$ on manifolds with bounded geometry, a result that was used for the analysis of the Stokes various classes of manifolds, including manifolds with cylindrical ends, in [63, 67, 68]. Manifolds with cylindrical ends appear in may applications that involve “wave-guides” or “layers” (see e.g. [19, 20, 21, 22, 23, 24, 25, 42, 63, 68, 74, 137]). Let us finally add that the main operator in the structure of the Stokes and Navier-Stokes equations on Riemannian manifolds is the deformation Laplacian $\boldsymbol{L}:=2\operatorname{Def}\,^{*}\operatorname{Def}\,$. In their groundbreaking paper [40], Ebin and Marsden described the deformation Laplacian as the right Laplace type operator to describe the Navier-Stokes equation on compact Riemannian manifolds. Pierfelice [108] also used the deformation-Laplacian for the definition and the analysis of the Navier-Stokes equation on a non-compact Riemannian manifold with negative Ricci curvature. This is the choice that we also adopt in the description of our Stokes operator $\boldsymbol{\Xi}$ of Equation (1.3). Our generalized Stokes operator is closely related to the Brinkman operator. The Brinkman flow was studied by Karageorghis, Lesnic, and Marin in [57, 58], and Bacuta, Hassell, Hsiao, and Sayas in [26]. A class of solutions to the Navier-Stokes equations for flows in plane-parallel channels and related singular perturbation problems were studied by Mazzucato and Taylor [86].
@@ 1.5. Acknowledgements
We thank Massimo Lanza de Cristoforis, Dorina, Irina, and Marius Mitrea, Sergey Mikhailov, and Mihai Putinar for useful discussions.
@@ 2.2. An abstract integration by parts formula
    or any vector field $X$, we have [107, page 49]
@@ 3.4. Manifolds with cylindrical ends, compatible vector bundles and differential operators
in which case ${\mathcal{R}_{\infty}}\nabla^{LC}=\nabla^{LC}$.
@@ 4.1. General results on normal lateral limits of pseudodifferential operators
The result follows by applying Lebesgue’s dominated convergence theorem. See [41, Appendix E]. ∎
@@ 5. Pseudodifferential operators on manifolds with straight cylindrical ends
From now on, $M$ will be a manifold with straight cylindrical ends, $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$. Also, in the following, $E\to M$ will be a fixed compatible Hermitian vector bundle. We now define two needed classes of pseudodifferential operators on $M$, the $\operatorname{inv}$-calculus” and the $\operatorname{ess}$-calculus.” For $M$ closed, these definitions are trivial (they recover the usual pseudodifferential operators, so the give nothing new), so the only interesting case is the one when $M$ is not compact. Including the case $M$ compact is useful, however, in order to formulate our results in a more general setting, as they are true for both closed manifolds and for manifolds with straight cylindrical ends.
The resulting classes of pseudodifferential operators are very closely related to the $b$- and $c$- calculi of Melrose [93, 84]. Related calculi have also been considered by Schulze and Schrohe [121, 122, 124]. See [67] for more references and for a discussion of the relation between our calculi and the ones of Mazzeo, Melrose, Schrohe, and Schulze. The results of this section are close to those in [63, 67]. The presentation is, however, different and, occasionally, it provides more general results. For instance, in [63], we work only with classical pseudodifferential operators, whereas here we allow the Kohn-Nirenberg classes of symbols recalled in Definition 3.5.
@@ 5.5. Limit operators
    If $M=M^{\prime}\times\mathbb{R}$ and $P$ is translation invariant (that is, $P\in\Psi_{\operatorname{inv}}^{m}(M^{\prime}\times\mathbb{R};E,F)^{\mathbb{R}}$), then ${\color[rgb]{1,0,1}{\mathcal{R}_{\infty}}(P)}=P$.
@@
${\color[rgb]{1,0,1}{\mathcal{R}_{\infty}}(P)}\in\psi_{\operatorname{sus}}^{m}(M^{\prime};E,F)\mathrel{\,:=\,}\Psi^{m}(M\times\mathbb{R};E,F)^{\mathbb{R}}$
be as in Lemma 5.11. Then $\color[rgb]{1,0,1}{\mathcal{R}_{\infty}}(P)$} is called the limit operator associated to $P$. The resulting map
${\mathcal{R}_{\infty}}:\Psi_{\operatorname{inv}}^{m}(M;E,F)\ni P\longmapsto{\color[rgb]{1,0,1}{\mathcal{R}_{\infty}}(P)}\in\psi_{\operatorname{sus}}^{m}(M^{\prime};E,F)$
@@ 5.6. The structure of the $\operatorname{inv}$-calculus
We also have the following characterization of Sobolev spaces [67], [63].
@@ 5.9. Translation invariance, the Fourier transform, and indicial operators
${\color[rgb]{1,0,1}{\mathcal{R}_{\infty}}(P)}(u)\mathrel{\,:=\,}\lim_{s\to\infty}\Phi_{-s}\big[P\Phi_{s}(u)|_{\{t<0\}}\big]\,.$
@@
6. Normal lateral limits on manifolds with straight cylindrical ends
In this section, we adapt the results from Section 4 on normal lateral limits on half-spaces to smooth, open domains in manifolds with cylindrical ends. We thus continue to assume that $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ is a smooth Riemannian manifold with straight cylindrical ends (Definition 3.11) and that $E,F\to M$ are two compatible vector bundle (see Definition 3.13). Most importantly, we let $\Omega\subset M$ be an open subset with smooth boundary $\Gamma:=\partial\Omega\neq\emptyset$ (so the assumption is that $\Gamma$ is also a smooth manifold) that is compatible with the structure of cylindrical ends of $M$, as in the introduction, Equation (1.2). That is:
@@ 6.1. Normal tubular neighborhoods
It is well-known that if $\Gamma$ is compact or if $M$ has straight cylindrical ends and $\Omega$ is compatible with this structure, then $\Gamma$ will have an $\epsilon$-normal tubular neighborhood, for some $\epsilon>0$ small enough, see [107, Corollary 5.5.3]. The curves $t\mapsto\exp(t\boldsymbol{\nu}(x))$, $x\in\partial\Omega$, will be called the normal geodesics to $\partial\Omega$. If $u$ is a section of $E$ over $M$, $\Gamma$ has an $\epsilon$-normal tubular neighborhood, and $t\in(-\epsilon,\epsilon)$, we let
@@ 6.2. Normal lateral limits for manifolds with cylindrical ends
Recall from Notation 6.2 that $\boldsymbol{\nu}$ is a fixed vector field on $M$ that is the outer unit normal vector to $\Gamma:=\partial\Omega$. Also, recall that $\sharp:TM\to T^{*}M$ is the isomorphism defined by the metric. We now turn to the study of the normal lateral limits of operators of order m=1m=-1. We first need to establish some more notation.
@@ 6.3. Douglis-Nirenberg operators: jumps, ellipticity, and Fredholmness
The given formula for principal symbol of $\boldsymbol{\Xi}$ is a consequence of [130, Chapter 5 §12, (12.20)] (but see also Corollary 2.6 for the principal symbol of $\operatorname{Def}^{*}\,\operatorname{Def}\,$ and Remark 2.8(iii) for the principal symbol of $\nabla$). The result then follows by verifying that the product of these two matrices is the identity. ∎
See also [68, Theorem 5.11 and Remark 5.12] and [36]. We notice that the formula for the inverse of the principal symbol was obtained using Schur’s complement formula.
@@
We will use the Douglis-Nirenberg-ellipticity of $\boldsymbol{\Xi}$ to establish (in the next subsection, under certain positivity assumptions on the potentials $V$ and $V_{0}$) its Fredholm property. To that end, we shall use the following characterization of the Fredholm property in the essentially translation invariant calculus. (This result was discussed and proved in great detail in [63, Theorem 15.4.19] and [67, Theorem 7.12], so we will not include a proof here. It builds on earlier results of many people, including Kondratiev, Melrose, and Schulze, see the above quoted works for references.) Recall the notation of the last remark, Remark 6.14. Then we have the following characterization of Fredholm property for a Douglis-Nirenberg-elliptic operator on a manifold with straight cylindrical ends (see [68] and [63]).
@@ 7. Green formulas and invertibility of $\boldsymbol{\Xi}$ on cylinders
We thus assume in this section that $M:=\mathfrak{S}\times\mathbb{R}$, where $\mathfrak{S}$ is a closed manifold (i.e., smooth, compact, without boundary). Moreover, all other objects used in this section will be translation invariant, thus $\Omega=\Omega^{\prime}\times\mathbb{R}$ and $V$ and $V_{0}$ descend to functions on $\mathfrak{S}$, denoted, respectively, $\color[rgb]{1,0,1}\widehat{V}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}$ (see Equation (5.6) for a justification of this notation). Explicitly, we have $V(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}(x^{\prime})$ and $V_{0}(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}_{0}(x^{\prime})$ for $x^{\prime}\in\mathfrak{S}$ and $t\in\mathbb{R}$. The results of this section will be used for $\mathfrak{S}$ possibly disconnected, so we will not assume that $\mathfrak{S}$ is connected, although the proofs of the results in this section can be reduced immediately to the connected case. Note also that we are continuing to assume that $M$ is a manifold with cylindrical ends, albeit of a very particular kind. Recall that the scalar product on $L^{2}(M;E)$ is

### 28 May 2026 v2

@@ 1.4. Overview of the existing literature
The analysis on non-compact manifolds poses more challenges than that the one on compact manifolds. A significant issue in this sense is the construction of the classes of pseudodifferential operators. In this regard, we refer to the contributions of Ammann, Lauter and Nistor [13, 14], Mazzeo and Melrose [105], Melrose [114], Melrose and Mendoza [116], Nistor [125], Nistor, Weinstein, and Xu [126] Schrohe [143, 144], Schulze [145, 146], Schulze and Wong [147], Shubin [151], to mention just a few among very many others. For instance, some results relevant for boundary value problems were obtained by Grieser [58, 63], who used the Calderon projection in the framework of the cusp-edge calculus [105] for related boundary value problems for (classically) elliptic operators. The cusp-edge calculus is a completion of a groupoid calculus [126] corresponding to a structural Lie algebra of vector fields that is, in fact, more general than the one appearing in our setting of manifolds with straight cylindrical ends.
Another important issue in the analysis on non-compact manifolds is finding the “right definition” of the relevant function spaces. In this paper, as in [86], we have defined our function spaces by using connections. See also [5, 6, 142]. Amann [6] and Ammann, Große, Nistor [8, 10] studied function spaces on manifolds with bounded geometry and obtained well-posedness results for the Laplace operator. Using the uniform Shapiro-Lopatinski conditions, Große and Nistor [65] obtained regularity results for elliptic boundary value problems on manifolds with bounded geometry. Lie manifolds are special manifolds with bounded geometry [12, 13], which allow the implementation of the layer potential methods.
@@
In [64], Große, Kohr, and Nistor have obtained the $L^{2}$-unique continuation property for the deformation operator $\operatorname{Def}\,$ on manifolds with bounded geometry, a result that was used for the analysis of the Stokes various classes of manifolds, including manifolds with cylindrical ends, in [83, 87, 88]. Manifolds with cylindrical ends appear in may applications that involve “wave-guides” or “layers” (see e.g. [23, 24, 25, 26, 27, 28, 29, 30, 52, 54, 56, 83, 88, 94, 161]). Let us finally add that the main operator in the structure of the Stokes and Navier-Stokes equations on Riemannian manifolds is the deformation Laplacian $\boldsymbol{L}:=2\operatorname{Def}\,^{*}\operatorname{Def}\,$. In their groundbreaking paper [50], Ebin and Marsden described the deformation Laplacian as the right Laplace type operator to describe the Navier-Stokes equation on compact Riemannian manifolds. Pierfelice [131] also used the deformation-Laplacian for the definition and the analysis of the Navier-Stokes equation on a non-compact Riemannian manifold with negative Ricci curvature. This is the choice that we also adopt in the description of our Stokes operator $\boldsymbol{\Xi}$ of Equation (1.4). Our generalized Stokes operator is closely related to the Brinkman operator. The Brinkman flow was studied by Karageorghis, Lesnic, and Marin in [77, 78], and Băcuţă, Hassell, Hsiao, and Sayas in [31]. A class of solutions to the Navier-Stokes equations for flows in plane-parallel channels and related singular perturbation problems were studied by Mazzucato and Taylor [107].
@@ 1.5. Acknowledgements
We thank Massimo Lanza de Cristoforis, Dorina, Irina, and Marius Mitrea, Sergey Mikhailov, and Mihai Putinar for useful discussions. We also thank Daniel Grieser and Karl-Mikael Perfekt for useful suggestions.
@@ 2.2. An abstract integration by parts formula
    or any vector field $X$, we have [130, page 49]
@@ 3.4. Manifolds with cylindrical ends, compatible vector bundles and differential operators
in which case ${\mathcal{R}_{\infty}}\nabla^{LC}=\nabla^{LC}$. The choice of manifolds (i.e., with straight cylindrical ends) and of the associated functions and differential operators is motivated by practical questions, since these are the objects that are the most likely to appear in practical applications. However, our results extend to the more general case of manifolds with cylindrical ends (not necessarily straigh, but asymptotically straight). One just needs to slightly enlarge the class of functions, vector bundles, differential operators, and, especially, pseudodifferential operators. These questions will be considered in another paper.
@@ 4.1. General results on normal lateral limits of pseudodifferential operators
The result follows by applying Lebesgue’s dominated convergence theorem. See [55, Appendix E]. ∎
@@ 5. Pseudodifferential operators on manifolds with straight cylindrical ends
From now on, $M$ will be a manifold with straight cylindrical ends, $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$. Also, in the following, $E\to M$ will be a fixed compatible Hermitian vector bundle. We now define two needed classes of pseudodifferential operators on $M$, the $\operatorname{inv}$-calculus” and the $\operatorname{ess}$-calculus.” Our presentation is complete, but short (unlike the previous sections). The reason is that there are many works on related classes of pseudodifferential operators, beginning with the $b$- and $c$- calculi of Melrose [114, 105]. Related calculi have also been considered by Schulze and Schrohe [144, 145, 147]. See [87] for more references and for a discussion of the relation between our calculi and the ones of Mazzeo, Melrose, Schrohe, and Schulze. See also [62, 100, 101, 117, 147].
The results of this section are close to those in [83, 87], but the presentation is different and, occasionally, it provides more general results. For instance, in [83], we work only with classical pseudodifferential operators, whereas here we allow the Kohn-Nirenberg classes of symbols recalled in Definition 3.5.
@@ 5.5. Limit operators
    If $M=M^{\prime}\times\mathbb{R}$ and $P$ is translation invariant (that is, $P\in\Psi_{\operatorname{inv}}^{m}(M^{\prime}\times\mathbb{R};E,F)^{\mathbb{R}}$), then ${\mathcal{R}_{\infty}}(P)=P$.
@@
${\mathcal{R}_{\infty}}(P)\in\psi_{\operatorname{sus}}^{m}(M^{\prime};E,F)\mathrel{\,:=\,}\Psi^{m}(M\times\mathbb{R};E,F)^{\mathbb{R}}$
be as in Lemma 5.11. Then ${\mathcal{R}_{\infty}}(P)$ is called the limit operator associated to $P$. The resulting map
${\mathcal{R}_{\infty}}:\Psi_{\operatorname{inv}}^{m}(M;E,F)\ni P\longmapsto{\mathcal{R}_{\infty}}(P)\in\psi_{\operatorname{sus}}^{m}(M^{\prime};E,F)$
@@ 5.6. The structure of the $\operatorname{inv}$-calculus
We also have the following characterization of Sobolev spaces [87, 83].
@@ 5.9. Translation invariance, the Fourier transform, and indicial operators
${\mathcal{R}_{\infty}}(P)(u)\mathrel{\,:=\,}\lim_{s\to\infty}\Phi_{-s}\big[P\Phi_{s}(u)|_{\{t<0\}}\big]\,.$
@@
6. Main properties of the $\operatorname{inv}$ and $\operatorname{ess}$ calculi
In this section, we establish some of the main properties of the operators in the $\operatorname{inv}$” and $\operatorname{ess}$” calculi of the previous section. We concentrate on the properties needed in the study of layer potentials. We first adapt the results from Section 4 on normal lateral limits on half-spaces to smooth, open domains in manifolds with cylindrical ends and our calculi of the previous section ($\operatorname{inv}$ and $\operatorname{ess}$). Then we recall Douglis-Nirenberg-operators and their Fredholm, (elliptic) regularity, and spectral invariance properties.
We thus continue to assume that $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ is a smooth Riemannian manifold with straight cylindrical ends (Definition 3.11) and that $E,F\to M$ are two compatible vector bundle (see Definition 3.13). Most importantly, we let $\Omega\subset M$ be an open subset with smooth boundary $\Gamma:=\partial\Omega\neq\emptyset$ (so the assumption is that $\Gamma$ is also a smooth manifold) that is compatible with the structure of cylindrical ends of $M$, as in the introduction, Equation (1.3). That is:
@@ 6.1. Normal tubular neighborhoods
It is well-known that if $\Gamma$ is compact or if $M$ has straight cylindrical ends and $\Omega$ is compatible with this structure, then $\Gamma$ will have an $\epsilon$-normal tubular neighborhood, for some $\epsilon>0$ small enough, see [130, Corollary 5.5.3]. The curves $t\mapsto\exp(t\boldsymbol{\nu}(x))$, $x\in\partial\Omega$, will be called the normal geodesics to $\partial\Omega$. If $u$ is a section of $E$ over $M$, $\Gamma$ has an $\epsilon$-normal tubular neighborhood, and $t\in(-\epsilon,\epsilon)$, we let
@@ 6.2. Normal lateral limits for manifolds with cylindrical ends
By extending the above mapping property to Besov spaces, one can extend our well-posedness result to $L^{p}$-type Sobolev and Besov spaces. This will be discussed in a future publication. Recall from Notation 6.2 that $\boldsymbol{\nu}$ is a fixed vector field on $M$ that is the outer unit normal vector to $\Gamma:=\partial\Omega$. Also, recall that $\sharp:TM\to T^{*}M$ is the isomorphism defined by the metric. We now turn to the study of the normal lateral limits of operators of order m=1m=-1. We first need to establish some more notation.
@@ 6.3. Douglis-Nirenberg operators: jumps, ellipticity, and Fredholmness
The given formula for principal symbol of $\boldsymbol{\Xi}$ is a consequence of [154, Chapter 5 §12, (12.20)] (but see also Corollary 2.6 for the principal symbol of $\operatorname{Def}^{*}\,\operatorname{Def}\,$ and Remark 2.8(iii) for the principal symbol of $\nabla$). The result then follows by verifying that the product of these two matrices is the identity. ∎
See also [88, Theorem 5.11 and Remark 5.12] and [46]. We notice that the formula for the inverse of the principal symbol was obtained using Schur’s complement formula.
@@
We will use the Douglis-Nirenberg-ellipticity of $\boldsymbol{\Xi}$ to establish (in the next subsection, under certain positivity assumptions on the potentials $V$ and $V_{0}$) its Fredholm property. To that end, we shall use the following characterization of the Fredholm property in the essentially translation invariant calculus. (This result was discussed and proved in great detail in [83, Theorem 15.4.19] and [87, Theorem 7.12], so we will not include a proof here. It builds on earlier results of many people, including Kondratiev, Melrose, and Schulze, see the above quoted works for references.) Recall the notation of the last remark, Remark 6.14. Then we have the following characterization of Fredholm property for a Douglis-Nirenberg-elliptic operator on a manifold with straight cylindrical ends (see [88] and [83]).
@@
6.4. Further properties: regularity and spectral invariance
We shall repeatedly make use of the following spectral invariance and elliptic regularity results from [83, 87]. We have the following generalization of the classical elliptic regularity result (see Theorem 15.3.30 of [83] or [87, Proposition 6.5], see also [115, 145]). We use the notation of Definition 6.11. Thus $E_{j},F_{i}\to M$, $i,j=1,\ldots,k$, are compatible vector bundles, $E=\oplus_{j=1}^{k}E_{j}$, and $F=\oplus_{i=1}^{k}F_{i}$, and $\boldsymbol{s}=(s_{1},\ldots,s_{k})$ and $\boldsymbol{t}=(t_{1},\ldots,t_{k})$, $s_{i},t_{j}\in\mathbb{R}$.
Theorem 6.16.
Let $P\in\Psi_{\operatorname{ess}}^{[\boldsymbol{s}+\boldsymbol{t}]}(M;E,F)$ be Douglis-Nirenberg-elliptic. Let also $m,m^{\prime}\in\mathbb{R}$. If $u\in H^{m^{\prime}+[\boldsymbol{t}]}(M;E)$ is such that $Pu\in H^{m-[\boldsymbol{s}]}(M;F)$, then $u\in H^{m+[\boldsymbol{t}]}(M;E)$.
We have the following related spectral invariance property ([83, Theorem 15.4.11], see also [20, 87, 120]).
Theorem 6.17.
Let $s_{i},t_{j}\geq 0$, $m\in\mathbb{R}$, and $T\in\Psi_{\operatorname{ess}}^{[\boldsymbol{s}+\boldsymbol{t}]}(M;E,F)$. Let
$\mathcal{D}(T_{\rm{max}})\mathrel{\,:=\,}\{\,\xi\in H^{m-[\boldsymbol{s}]}(M;E):\ T\xi\in H^{m-[\boldsymbol{s}]}(M;F)\,\}$
and assume that $T:\mathcal{D}(T_{\rm{max}})\to H^{m-[\boldsymbol{s}]}(M;E)$ is Fredholm with pseudo-inverse denoted $T^{(-1)}$. If one of the $s_{i}$ or $t_{i}$ is (strictly) positive (i.e., >0>0), we also assume that $T$ is $(\boldsymbol{s},\boldsymbol{t})$-Douglis-Nirenberg-elliptic. Then its Moore-Penrose pseudo-inverse $T^{(-1)}\in\Psi_{\operatorname{ess}}^{[\mathbf{-s-t}]}(M;F,E)$.
This result was proved in case $T$ is actually invertible in the above mentioned publications, so let us include a proof for the statement in case $T$ is Fredholm, using the corresponding statement for “$T$ invertible”.
Proof.
We can assume that $E=F$ and $\boldsymbol{s}$ and $\boldsymbol{t}=0$ are constant, so we work with the usual order (this is the only case used in this paper). Let us notice that $T$ is elliptic (if it is of order zero, this follows from the Fredholm property, Theorem 6.15). Elliptic regularity, Theorem 6.16 gives that its kernel $\mathcal{N}$ consists of functions in $H^{\infty}(M)$ and hence the orthogonal projection onto $p_{\mathcal{N}}\in\Psi_{\operatorname{ess}}^{-\infty}(M;E)$. Then $p_{\mathcal{N}}+T^{*}T$ is invertible and hence $(p_{\mathcal{N}}+T^{*}T)^{-1}\in\Psi_{\operatorname{ess}}^{[-2\boldsymbol{s}]}(M;E)$. The desired result follows from $T^{(-1)}=(p_{\mathcal{N}}+T^{*}T)^{-1}T^{*}\in\Psi_{\operatorname{ess}}^{[-2\boldsymbol{s}]}(M;E)$. ∎
These results are true (and were known for a long time) if $M$ closed, in which case we replace $\Psi_{\operatorname{ess}}^{m}(M;E)$ with $\Psi^{m}(M;E)$.
@@ 7. Green formulas and invertibility of $\boldsymbol{\Xi}$ on cylinders
We thus assume in this section that $M:=\mathfrak{S}\times\mathbb{R}$, where $\mathfrak{S}$ is a closed manifold (i.e., smooth, compact, without boundary). Moreover, all other objects used in this section will be translation invariant, thus $\Omega=\Omega^{\prime}\times\mathbb{R}$ and $V$ and $V_{0}$ descend to functions on $\mathfrak{S}$, denoted, respectively, $\widehat{V}$ and $\widehat{V}_{0}$ (see Equation (5.6) for a justification of this notation). Explicitly, we have $V(x^{\prime},t)=\widehat{V}(x^{\prime})$ and $V_{0}(x^{\prime},t)=\widehat{V}_{0}(x^{\prime})$ for $x^{\prime}\in\mathfrak{S}$ and $t\in\mathbb{R}$. The results of this section will be used for $\mathfrak{S}$ possibly disconnected, so we will not assume that $\mathfrak{S}$ is connected, although the proofs of the results in this section can be reduced immediately to the connected case. Note also that we are continuing to assume that $M$ is a manifold with cylindrical ends, albeit of a very particular kind. Recall that the scalar product on $L^{2}(M;E)$ is